IR Demo Project Weekly Report for February 22-26, 1999
Highlights for the week include: achieving a lasing power record (553 W cw), recirculating up to 3.7mA cw electron beam while lasing, and lasing stably at 500+ W cw with ~3 mA recirculating beam (400+ W for hours is now "easy" to achieve).
Test Plans for the first user demonstration tests have been signed off for User Lab 1. The first test will be Michelle Shinn's spectroscopic measurements.
On Feb. 24th, the chair of the DOE-BES Panel on "Novel and Coherent Light Sources", Prof. Steve Leone (Univ. of CO) presented the panel report to the Basic Energy Sciences Advisory Committee (BESAC). The panel recommended that DOE should plan to spend the order of $6M a year for the next two years for R&D for the so-called "4th Generation Light Source", which is assumed to be a hard x-ray source based on a linac driven FEL. In comparison to the potential x-ray applications, more near term FEL applications in the IR and UV did not excite the panel. However, given the relevance of the present IR Demo to some of the technical issues that are important for 4th Generation Sources, JLab should be able to attract some of the potential R&D funds, particularly for Coherent Synchrotron Radiation (CSR) experiments, other accelerator physics issues pertaining to the transport of high brightness electron beams, and the important problem of synchronizing external lasers to FELs.
FEL Installation/Maintenance Activities
The machine was up and running continuously all week, save for recesiations of the cathode Monday and Wednesday evenings. Consequently there was not much in the way of installation and maintenance this week, although work continues on instrumenting Laser Safety Systems for the User Labs.
FEL Commissioning Activities
We have been running the FEL since 0800 Monday, 22 Feb. 99. The cathode has delivered ~240C total since 10 Feb. 99, the day we turned the gun on. At this writing (0935 on 26 Feb. 99) we have the FEL lasing stably at 400+ W with ~2.5 mA beam; we are letting the machine sit untouched and monitoring its stability. At the moment the cathode can "only" provide 2.5 mA at 74.8 MHz (~33 pC bunch charge); we will recesiate the cathode on Monday, on which Day Shift is devoted to machine maintenance. We had also recesiated the cathode successfully twice this week, on Monday and Wednesday evenings. Running at ~3 mA current, the cathode e-folding lifetime appears to be (very roughly!) 12 hours. On Wednesday we achieved 533 W cw power output at 4.9 Ám with ~3 mA average current. Yesterday we recirculated 3.7 mA stably, but tripped on a beam-loss monitor as we were optimizing the settings of the FEL optics. We reached 520 W yesterday. At this point we suspect, very tentatively in view of other details that are not yet well understood, that mirror heating may limit us from reaching substantially more power. ( We remind the audience that we are still running on so-called:" first light" optics which were designed for the 100-200 watt first light commissioning effort.) According to our planned evolution of optics , We have prepared a silicon mirror that would be more thermally tolerant, and we are debating the optimal time to install it.
The machine is behaving very well and generally seems to be reproducible. For example, yesterday we loaded a file of saved settings from the previous day's high-power run, and in minutes were lasing at 470 W with ~3 mA current. We are cataloging improvements to make during the next rebuild of the gun, such as improving the stability of the buncher cavity and improving the capability of a couple of the viewers.
At some point we may need to embark on a program of injector optimization to get very high
currents (~5 mA), but at this time the injector does not appear to be impeding our progress. We also need to decide on a machine-safe work-around of the trips in the interlocks of the srf cavity
waveguide IR detectors. We have been taking careful data during the high-power runs with this goal in mind, while invoking administrative control over the detectors. The interlocks are meant to prevent damage to the polyethylene warm windows in the cryomodule. The data suggests strongly that these windows are not being heated significantly during our runs.
Measurements of electron-beam quality continue. Normalized transverse emittance at the wiggler seems to be varying between 10 and 20 mm-mrad for reasons that are not yet fully understood. A very nice set of data on transverse emittance versus cryomodule gang phase (related to bunch length) was taken at the exit of the first recirculation arc. It suggests the presence of coherent synchrotron radiation effects, but it is not definitive. A carefully controlled parametric study remains to be done, and part of the program for diagnostics is to establish a good machine setup to support the study. The machine setup for CSR studies and the machine setup for high-power lasing are not necessarily one and the same.
This evening we will work toward commissioning the first "user test" in User Lab 1. The Laser
Safety System and safety documentation are now in order to support the test. We plan to devote
second shift activities to the user tests as we continue high power commissioning efforts during the first shifts next week.